### v2

# SFO: Piloting VLM Feedback for Offline RL

Abstract
While internet-scale image and textual data have enabled strong generalization in Vision-Language Models (VLMs), the absence of internet-scale control data has impeded the development of similar generalization in standard reinforcement learning (RL) agents. Although VLMs are fundamentally limited in their ability to solve control tasks due to their lack of action-conditioned training data, their capacity for image understanding allows them to provide valuable feedback in RL tasks by recognizing successful outcomes. A key challenge in Reinforcement Learning from AI Feedback (RLAIF) is determining how best to integrate VLM-derived signals into the learning process. We explore this question in the context of offline RL and introduce a class of methods called Sub-Trajectory Filtered Optimization (SFO). We identify three key insights. First, trajectory length plays a crucial role in offline RL, as full-trajectory preference learning exacerbates the stitching problem, necessitating the use of sub-trajectories. Second, even in Markovian environments, a non-Markovian reward signal from a sequence of images is required to assess trajectory improvement, as VLMs do not interpret control actions and must rely on visual cues over time. Third, a simple yet effective approach—filtered and weighted behavior cloning—consistently outperforms more complex RLHF-based methods. We propose Sub-Trajectory Filtered Behavior Cloning (SFBC), a method that leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures to improve robustness and prevent turbulence. This study is preliminary; we provide initial evidence through evaluations on a toy control domain. Please enjoy our airport puns.
1 Introduction
A long-standing goal in reinforcement learning (RL) is to develop agents capable of solving a diverse set of tasks (i.e., broad generalization) while also learning efficiently from limited experience (i.e., rapid adaptation). Despite substantial progress, RL remains highly sample-inefficient and slow to adapt to new environments. Addressing these inefficiencies requires a foundation model for RL—one that generalizes across tasks and improves learning efficiency. However, a major obstacle to achieving this is the absence of large-scale, high-quality control datasets.
Reinforcement Learning from AI Feedback (RLAIF) presents a promising approach by leveraging vision-language models (VLMs) as scalable sources of learning signals (Bai et al.,, 2022; Klissarov et al.,, 2023; Ma et al.,, 2023; Faldor et al.,, 2024; Rocamonde et al.,, 2024; Baumli et al.,, 2023; Lee et al.,, 2024; Wang et al.,, 2024). However, determining how best to integrate VLM-derived signals into RL remains an open challenge. A key difficulty is that VLMs struggle to discriminate between similar trajectories, making it difficult to provide precise, informative feedback. This challenge is further complicated in online RL, where agents must learn from randomly initialized policies, leading to highly ambiguous early-stage feedback.
Offline RL, in contrast, provides access to pre-collected data, which can include trajectories that are easier for VLMs to differentiate. For this reason, we focus on offline RL. However, integrating VLMs into offline RL introduces additional challenges, particularly in handling trajectory credit assignment and feedback propagation.
To address these issues, we introduce Sub-Trajectory Filtered Optimization (SFO), a class of methods that effectively incorporate VLM-derived feedback to reach new heights in offline RL. Our study identifies three key insights:
  1. 1.
    First, trajectory length is a crucial factor in offline RL—full-trajectory preference learning exacerbates the well-known stitching problem, necessitating the use of sub-trajectories, as depicted in Figure 1.
  2. 2.
    Second, even in Markovian environments, a non-Markovian reward model based on image sequences is required to assess trajectory improvement. Since VLMs are not trained on action-conditioned data, they have no basis for reasoning about how actions influence state transitions and must instead rely on visual cues over time. This is depicted in Figure 1. While discrete actions can sometimes be represented textually, many control tasks involve continuous actions (e.g., torques) that cannot be meaningfully encoded in a format suitable for VLMs. Furthermore, providing a reward at every time step by querying a VLM is computationally intractable for many problems without additional distillation Wang et al., (2024).
  3. 3.
    Third, despite the complexity of existing RLHF-based methods, we find that a simple yet effective approach—filtered and weighted behavior cloning—performs best in practice.
To that end, we introduce Sub-Trajectory Filtered Behavior Cloning (SFBC), which leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures. We empirically evaluate SFBC in a toy control domain, demonstrating its effectiveness compared to existing baselines, including naive behavior cloning and preference-based methods such as Direct Preference Optimization (DPO) (Rafailov et al.,, 2023). Notably, our method outperforms AWAC (Nair et al.,, 2020) and TD3+BC (Fujimoto and Gu,, 2021) when using ground-truth rewards, highlighting the effectiveness of VLM-driven sub-trajectory feedback.
(a) Comparing full trajectories can be uninformative, whereas evaluating sub-trajectories can resolve ambiguities and better capture meaningful distinctions.
(b) An example trajectory where the agent successfully swings up a pendulum, which should receive positive feedback.
(c) A failure case where Markovian feedback assigns the same reward to a bad trajectory as a good one, failing to differentiate between them.
Figure 1: Illustration of the necessity of Non-Markov sub-trajectory feedback. (a) Full-trajectory comparisons can be misleading, while sub-trajectories allow more informative distinctions. (b) A successful trajectory of a pendulum swing-up task. (c) A failure case where Markovian feedback cannot distinguish between successful and unsuccessful trajectories.
2 Related Work
One approach to improving RL efficiency is the manual collection of large-scale offline datasets (Vuong et al.,, 2023). However, this is difficult to scale due to the expense of generating diverse and high-quality trajectories on robotics in the real world. Similarly, human annotation of RL datasets is costly and difficult to scale, given the need for expert-level feedback and the challenge of obtaining consistent reward signals across tasks.
VLMs offer a promising alternative to human annotation by providing AI-generated feedback. Prior work has explored using VLMs for reward design and environment generation in code. Eureka and OMNI-EPIC (Ma et al.,, 2023; Faldor et al.,, 2024) have proposed methods where VLMs write programs to specify rewards and structure environments for learning. However, just as machine learning is needed to specify object categories (e.g., recognizing a dishwasher) rather than relying on hand-crafted rules, RL also requires machine-learned reward models rather than fixed rules from a VLM (e.g., to identify when a dishwasher has been unloaded). We firmly believe that hardcoded rules from a VLM cannot specify sufficiently complex tasks to elicit sufficiently robust reinforcement learning agents.
Recent work has also attempted to leverage AI-based reward detection to improve RLHF. Agent Q (Putta et al.,, 2024) uses AI-driven reward and success detection to generate preference labels for DPO. RL-VLM-F (Wang et al.,, 2024) distills a reward model from AI-generated labels but operates on Markovian states, which we find to be suboptimal. Rocamonde et al., (2024) and Baumli et al., (2023) use a CLIP model to provide Markovian feedback. Additionally, RoboCLIP (Sontakke et al.,, 2023) encodes text and video representations using CLIP to provide feedback. While RoboCLIP-style methods offer a useful way to encode feedback, we propose an end-to-end approach using a VLM (GPT-4o in our case), which provides greater flexibility. Our study focuses on how best to utilize this feedback and how to leverage feedback in an offline context.
Complementary approaches involve automatic environment generation and world modeling (Bruce et al.,, 2024; Valevski et al.,, 2024). These efforts provide a scalable way to generate diverse training environments, but they still require specifying reward signals to learn effectively. Similarly, automated task selection (Wang et al.,, 2023) and task ordering techniques such as Prioritized Level Replay (PLR) contribute to efficient learning but do not replace the need for automated high-quality feedback signals.
3 Methods
Our objective is to learn an optimal policy $\pi^{*}$ in an offline RL setting using vision-language model (VLM) feedback. Given a dataset $\mathcal{D}=\{(s_{t},a_{t},s_{t+1},a_{t+1},\dots)\}$ of trajectories collected by unknown behavior policies, we seek to leverage VLMs to extract meaningful rewards and preferences over trajectory segments while mitigating the challenges of stitching with offline RL and Markovian rewards with VLMs, depicted in Figure 1.
3.1 Sub-Trajectory Filtered Behavior Cloning (SFBC)
To incorporate VLM feedback effectively, we propose Sub-Trajectory Filtered Behavior Cloning (SFBC), which refines behavior cloning by selectively weighting and filtering trajectory segments based on VLM-derived success probabilities.
1. Sub-Trajectory Decomposition: Each trajectory $\tau=(s_{1},a_{1},s_{2},a_{2},\dots,s_{T})$ is divided into $n$ equal-length sub-trajectories $\tau_{i}$, where:
$\tau_{i}=(s_{i\cdot k},a_{i\cdot k},s_{i\cdot k+1},a_{i\cdot k+1},\dots,s_{(i+1)\cdot k}),$ (1)
with segment length $k=\lfloor T/n\rfloor$. Additionally, we subsample observations within each segment to reduce token usage when querying the VLM.
2. VLM-Based Filtering: A VLM is queried with both a Markov prompt and a non-Markov prompt. For example, in the Pendulum-v1 environment, we use:
  • Markov: “You are watching a video of a red stick. If the black dot is at the bottom of the stick, answer ‘Y’. Otherwise, answer ‘N’.”
  • Non-Markov: “You are watching a video of a red stick. If the stick has moved between sides of the screen (left to right or right to left), answer ‘Y’. Otherwise, answer ’N’.”
Each prompt is followed by a sequence of images representing the sub-trajectory. The probabilities of responses (“y”) and (“yes”) (after lower-casing the response) are summed, as are the probabilities of (“n”) and (“no”). We qualitatively find that the collective “no” probability is more reliable, and both the Markov and non-Markov prompts contribute essential information. Thus, the final success probability is computed as:
[…]
$P_{VLM}(\tau_{i})$ $=\min\left(1,P_{\text{Markov}}(\tau_{i})+P_{\text{Non-Markov}}(\tau_{i})\right).$ (4)
A sub-trajectory is retained if $P_{VLM}(\tau_{i})\geq\alpha$. To ensure robustness, we employ retrospective filtering, where a sub-trajectory $\tau_{i}$ is only included if the next sub-trajectory $\tau_{i+1}$ also meets the threshold:
$\mathcal{D}_{SFBC}=\{(s_{t},a_{t})\mid\tau_{i}\in\mathcal{D},P_{VLM}(\tau_{i})\geq\alpha\text{ and }P_{VLM}(\tau_{i+1})\geq\alpha\}.$ (5)
This filtering mechanism assumes that a failed sub-trajectory likely resulted from preceding failures, ensuring that faulty transitions are not reinforced.
3. Weighted Behavior Cloning: The learned policy, $\pi_{\theta}$, is then trained using the following behavior cloning objective:
$\mathcal{L}_{SFBC}=-\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}_{SFBC}}\left[P_{VLM}(\tau_{t})\log\pi_{\theta}(a_{t}|s_{t})\right].$ (6)
3.2 Comparison to Alternative SFO Methods
We contrast SFBC with alternative methods leveraging sub-trajectories, including:
  • VLM+TD3+BC: Interprets $P_{VLM}(\tau)$ as a reward and applies TD3+BC (Fujimoto and Gu,, 2021), a competitive offline RL algorithm.
  • S-DPO: Applies Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) to sub-trajectories using VLM-derived rankings.
4 Experiments
To evaluate the effectiveness of SFBC, we conduct experiments in the Pendulum-v1 environment, a standard continuous control benchmark. Our evaluation focuses on assessing SFBC’s ability to correctly identify and stitch together the most optimal segments from sub-optimal demonstrations while outperforming conventional offline RL baselines. While the VLM processes image-based observations to provide feedback, the learned policy itself conditions on lower-dimensional vector representations of the state, following standard RL formulations.
We compare SFBC against both standard offline RL algorithms and alternative VLM-assisted methods. To ensure robustness and statistical significance, we report success rate, mean return, and standard error across 15 seeds. Our analysis covers both overall performance comparisons and detailed ablation studies to isolate the contributions of each design choice in SFBC.
4.1 Dataset Construction
To evaluate the effectiveness of our proposed method in offline reinforcement learning, we construct a dataset for the Pendulum-v1 environment designed to explicitly test the ability to stitch optimal sub-trajectories together. Our dataset consists of 500 trajectories, each with a length of 600 time steps. These trajectories are generated from two distinct policies:
  • Expert Policy: A custom proportional-derivative (PD) controller designed to stabilize the pendulum in an upright position. The controller balances the pendulum when near vertical and applies a predefined acceleration-based heuristic otherwise.
  • Failure Policy: A manually constructed failure policy that consistently directs the pendulum downward, ensuring suboptimal performance.
To create a dataset that challenges the ability to stitch the best segments of given trajectories, each trajectory consists of one expert policy demonstration of length 300 and one failure policy demonstration of length 300, in a random order.
4.2 Baseline Comparisons
We compare our proposed method, SFBC, against a set of competitive baselines, including both standard offline RL algorithms and behavior cloning methods. Specifically, we evaluate:
  • BC Naive: A standard behavior cloning baseline trained on the entire dataset, including both expert and failure trajectories, without any filtering or weighting mechanisms.
  • VLM BC (Full-Trajectory): A VLM-assisted behavior cloning approach that applies success filtering at the full-trajectory level rather than sub-trajectories. Since our dataset contains a mix of expert and failure trajectories in random order, filtering at the full-trajectory level does not provide meaningful differentiation. We include this baseline primarily for didactic purposes.
  • TD3+BC (GT): A strong offline RL baseline using TD3+BC (Fujimoto and Gu,, 2021), trained with access to ground-truth rewards.
  • AWAC (GT): Another standard offline RL algorithm, AWAC (Nair et al.,, 2020), trained on the dataset with ground-truth rewards.
The results, presented in Table 1, demonstrate that SFBC outperforms all baselines in both mean return and success rate. Notably, SFBC achieves better performance than TD3+BC trained on ground-truth rewards, underscoring the effectiveness of leveraging VLM-derived trajectory filtering and weighting. This suggests that VLM-derived feedback captures non-myopic task success, effectively bypassing the need for explicit reward propagation, a known challenge in offline RL due to overestimation biases. Additionally, SFBC surpasses S-DPO. We hypothesize that S-DPO fails due to inherent limitations in existing vision-language models, particularly their difficulty in reasoning over multiple sequences of visual data, which is essential for accurate trajectory ranking.
Furthermore, SFBC outperforms VLM+TD3+BC, justifying our choice to treat $P_{VLM}$ as a weight rather than a reward. Conceptually, the weighting scheme can be interpreted as a value function $Q^{\pi_{\theta}}$, such that the weighted loss, $\mathcal{L}_{SFBC}$, implicitly optimizes a policy gradient. However, $\mathcal{L}_{SFBC}$ lacks an importance sampling correction, and $P_{VLM}$ is computed from offline data rather than the current policy, making it a better model of $Q^{\pi_{\text{behavior}}}$ (or $Q^{\pi_{\text{expert}}}$ after filtering). SFBC also resembles AWAC (Nair et al.,, 2020), except that the weighting $P_{VLM}$ should likewise be interpreted as approximating $e^{A^{\pi_{\text{expert}}}}$, rather than the advantage of the learned policy $e^{A^{\pi_{\theta}}}$. While SFBC does not fully correspond to a policy gradient, our results show that avoiding a direct reward interpretation—and the instability it introduces in reward propagation—is beneficial in practice.
Method Success Rate (%) Std. Error (%) Mean Return Std. Error
BC Naive 33 12 -4716 790
TD3+BC (GT) 27 11 -5131 814
VLM BC (Full-Trajectory) 13 9 -5234 578
AWAC (GT) 0 0 -7840 308
SF-BC (Ours) 73 11 -1585 518
VLM+TD3+BC 27 11 -5013 649
S-DPO 0 0 -6859 181
Table 1: Performance comparison of SFBC against standard offline RL and VLM-assisted baselines in the Pendulum-v1 environment. SFBC achieves the highest success rate and mean return, demonstrating its effectiveness in leveraging VLM feedback for sub-trajectory optimization.
4.3 Ablation Studies
We conduct ablation studies to assess the necessity of each component in SFBC. The results, presented in Table 2, confirm that each design choice contributes significantly to performance:
  • No Filtering: Removing the filtering mechanism results in significantly worse performance, demonstrating the necessity of discarding low-confidence sub-trajectories.
  • Markov Prompt Only: Using only the Markov prompt leads to degraded performance, confirming the importance of a non-Markovian understanding of trajectory improvement, even in Markovian environments, due to the inability of VLMs to process control data.
  • No Weighting: Without weighting, the method lacks prioritization of high-quality sub-trajectories, leading to reduced performance.
  • No Retrospective Filtering: This ablation resulted in the largest performance degradation. Without retrospective filtering, failure states remain in the dataset, potentially reinforcing poor decisions and leading to compounding errors in deployed policies. This highlights the importance of discarding preceding sub-trajectories that contribute to unsuccessful outcomes.
The ablations validate our hypothesis that all proposed modifications are necessary, further reinforcing the importance of VLM-informed filtering and weighting in offline RL.
Ablation Success Rate (%) Std. Error (%) Mean Return Std. Error
SF-BC (Ours) 73 11 -1585 518
No Filtering 40 13 -4164 883
Markov Prompt Only 40 13 -4229 869
No Weighting 33 12 -3459 604
No Retrospective Filtering 13 9 -5562 525
Table 2: Ablation study on SFBC, evaluating the impact of key design choices. Removing filtering, non-Markovian feedback, weighting, or retrospective filtering significantly degrades performance, demonstrating the necessity of each component.
5 Conclusion
In this work, we introduced Sub-Trajectory Filtered Behavior Cloning (SFBC), a method that leverages vision-language model (VLM) feedback for offline reinforcement learning by selectively filtering and weighting sub-trajectories. Our results demonstrate that SFBC outperforms standard offline RL methods such as TD3+BC and AWAC, as well as VLM-assisted baselines, validating the effectiveness of leveraging VLM-derived success probabilities for policy learning.
Our study provides key insights into the integration of VLM feedback with offline RL. First, sub-trajectory selection is critical, as full-trajectory preference learning exacerbates the stitching problem, making effective credit assignment impossible. Using sub-trajectories allows for improved learning stability and the ability to extract high-quality behaviors from mixed datasets. Second, non-Markovian feedback is essential, even in Markovian environments, since VLMs lack an understanding of control dynamics. Instead, they require vision-based trajectory analysis to correctly assess improvement over time. Our results confirm that non-Markovian prompts lead to significantly better performance. Finally, SFBC consistently outperforms competitive baselines, demonstrating that treating VLM-derived success probabilities as weights rather than rewards leads to superior offline policy learning. Our retrospective filtering mechanism further enhances performance by removing failure-propagating sub-trajectories.
Our findings suggest that VLM feedback provides non-myopic guidance, mitigating the need for explicit reward propagation and addressing key challenges in scaling offline RL. Most immediately, this work could be applied to automate feedback across a sufficiently diverse set of tasks to support training an RL foundation model from offline data. Such an application of SFBC would inherently fall under the category of offline meta-RL Beck et al., (2023), while also obviating the need for complex meta-gradient computation Vuorio et al., (2021). Putting such guidance systems on autopilot may finally provide the runway needed to arrive at a truly generalist agent for RL.
Another avenue for future research is exploring how offline feedback methods like SFBC can be extended to online RL. Since offline algorithms can be iteratively applied in an online setting, they provide a natural foundation for developing online reinforcement learning frameworks. Finally, while we leverage a vision-language model (VLM) trained primarily for static images, future work could explore the use of models specifically trained for video understanding, which may provide more accurate and temporally consistent trajectory assessments. As open-source and state-of-the-art VLMs continue to improve in video understanding, their ability to assess trajectory quality will become increasingly crucial for reinforcement learning applications to take off.
Appendix A Additional Experiments and Hyperparameters
A.1 Additional VLM+TD3+BC Experiments
We tested a variation where filtering was applied before TD3+BC, but it did not significantly improve performance. We also tested VLM+TD3+BC with the “Markov prompt only” to ensure rewards for TD3 were Markovian, as well as with a discount factor of $\gamma=0.5$ to encourage myopic behavior. Neither approach led to improved performance. The results are summarized in Table 3.
Method Return Mean Std Err Success Rate (%)
VLM+TD3+BC (Standard) -5013 649 27
VLM+TD3+BC (Filtered Before TD3+BC) -4576 660 27
VLM+TD3+BC (Markov Prompt Only) -4727 705 27
VLM+TD3+BC ($\gamma=0.5$) -5048 523 20
Table 3: Performance comparison of VLM+TD3+BC variants. Filtering before TD3+BC, using a Markov-only prompt, or reducing $\gamma$ did not significantly improve performance.
A.2 Prompting Details
For DPO, we used the following prompt:
“You are watching two videos of a red stick. The goal is to swing the stick up to gain height. It is bad to let it fall and lose momentum. Respond ’1’ if Video 1 is better, ’2’ if Video 2 is better.”
Followed by:
“Video 1:” $<$sub-trajectory$>$
“Video 2:” $<$sub-trajectory$>$
A.3 Hyperparameters
We used the default hyperparameters from D3RLpy for all baselines. Additionally, we tested a learning rate of $1e^{-3}$ for AWAC and TD3+BC, as this is the default for BC, but it did not improve results. We found $\alpha=0.1$ to be the most effective threshold for filtering.
For sub-trajectory processing, we used:
  • Sub-trajectory length: $100$
  • Subsampling factor: $20$
Success rates, mean returns, and standard errors are reported across 15 seeds to ensure statistical robustness.